# Mirror Online Conformal Prediction with Intermittent Feedback

### v3

@@ Abstract
Online conformal prediction enables the runtime calibration of a pre-trained artificial intelligence model using feedback on its performance. Calibration is achieved through set predictions that are updated via online rules so as to ensure long-term coverage guarantees. While recent research has demonstrated the benefits of incorporating prior knowledge into the calibration process, this has come at the cost of replacing coverage guarantees with less tangible regret guarantees based on the quantile loss. This work introduces intermittent mirror online conformal prediction (IM-OCP), a novel runtime calibration framework that integrates prior knowledge, while maintaining long-term coverage and achieving sublinear regret. IM-OCP features closed-form updates with minimal memory complexity, and is designed to operate under potentially intermittent feedback.
@@ I Introduction
As summarized in Table I, existing OCP schemes either cannot incorporate prior information or sacrifice long-term coverage guarantees (see Sec. II for details). To address these limitations, this work introduces intermittent mirror OCP (IM-OCP), a novel online calibration framework, that integrates prior knowledge, while ensuring long-term deterministic coverage guarantee and exhibiting constant memory and low-complexity updates like ACI. IM-OCP builds on online mirror descent (OMD) [21, 22, 23], inheriting from OMD also sublinear regret with respect to the quantile loss. We showcase the merits of IM-OCP for the task of received signal strength indicator (RSSI)-based localization under intermittent feedback [24, 25, 26].
@@ II-A Online Calibration via Set Prediction
We consider the problem of online calibration in sequential decision-making problems. At each round $t$, a learner observes an input $X_{t}\in\mathcal{X}$ and aims to predict the corresponding label $Y_{t}\in\mathcal{Y}$. To this end, the predictor assigns a negatively oriented score $s(X_{t},Y)$ to each value $Y\in\mathcal{Y}$, which is assumed to be upper bounded by a constant $B<\infty$, i.e., $s(X,Y)\leq B$. For example, the quadratic loss is defined as $s(X_{t},Y)=|f(X_{t})-Y|^{2}$, where $f(X_{t})$ is the prediction. Using predictor’s output, a prediction set $\mathcal{C}_{t}$ is produced that includes all labels $Y\in\mathcal{Y}$ with a score no larger than a threshold $r_{t}$, i.e., [10]
@@
In (2), the parameters $A$ and $\gamma$ are constants independent of $T$. By (2), the rate of miscoverage errors, $\mathrm{MisCov}(T)$, on an arbitrary sequence $\{(X_{t},Y_{t})\}_{t=1}^{T}$ converges to the target miscoverage level $\alpha$ as the number of steps $T$ tends to infinity.
@@ II-C Bayesian Adaptive Conformal Inference
In Appendix -A, we show that IB-ACI achieves sublinear expected regret (see (14)).
@@ III-A1 Regularizer Encoding Prior Information
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$ for any fixed $\sigma>0$. In Appendix -B, we show that the function $R(\cdot)$ is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\mu$.
@@ III-A2 Threshold Update via OMD
Using the regularizer $R(r)$ in (9), IM-OCP updates the threshold at each time $t$ according to the OMD rule [21, 22, 23]. To describe it, define the mirror map induced by function $R(r)$ as $\mathcal{M}_{R}(\cdot)=\nabla R(\cdot)$, and the corresponding inverse as $\mathcal{M}_{R}^{-1}(\cdot)=(\nabla R)^{-1}(\cdot)$. This mirror map $\mathcal{M}_{R}$ captures the geometry induced by the prior information. The update rule of IM-OCP is then given as
@@
When the feedback $E_{t-1}$ is observed ($\text{obs}_{t-1}$$=$$1$), this update corresponds to a gradient step in the dual space defined by function $R(\cdot)$. This adjusts the threshold $r_{t}$ in a direction informed by both the observed feedback and the prior knowledge embedded in the mirror map $\mathcal{M}_{R}$. In contrast, when no feedback is observed ($\text{obs}_{t-1}$$=$00), the update in (10) reduces to the identity map, and the threshold remains unchanged.
@@ V Conclusion
-A Regret Bound of IB-ACI
We now show that IB-ACI with a decreasing step size $\eta_{t}=t^{-1/2}/2$ enjoys a sublinear expected regret. Based on [22, Lemma 2.3], we have
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (16)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\left(\ell_{1-\alpha}(r_{i},r_{i}^{*})-\ell_{1-\alpha}(r_{i+1},r_{i}^{*})\right)\frac{\text{obs}_{i}}{p_{i}}.$
Noticing that $\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ is $\frac{\text{obs}_{i}}{p_{i}}$-Lipschitz, and with the additional assumption that $\psi(r)$ is $\epsilon$-Lipschitz, it follows that $h_{t}\psi(u)$ is $h_{t}\epsilon$-Lipschitz and that
$\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\ell_{1-\alpha}(r_{i+1},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (17)
$\leq\frac{\text{obs}_{i}}{p_{i}}|r_{i}-r_{i+1}|\leq\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}$
By combining (16) and (17), we obtain
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (18)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}$
Taking expectation over the sequence $\{\text{obs}_{i}\}^{T}_{i=1}$
$\mathrm{Reg}(T)\leq\mathbb{E}_{\text{obs}_{i}\sim p_{i}}\left[h_{t}\psi(u)+\sum_{i=1}^{T}\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}\right]$ (19)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{i}h_{i}\epsilon}\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}h_{i}\epsilon},$
where $p_{\text{min}}=\min\{p_{i}\}$. For the step size choice $\eta=1/(2\sqrt{t})$, we have $h_{t}=\mathcal{O}(\sqrt{t})$ and $1/h_{t}=\mathcal{O}(1/\sqrt{t})$. Thus, IB-ACI achieves sublinear regret
$\mathrm{Reg}(T)$ $\leq\mathcal{O}\left(\psi(u)\sqrt{T}+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}\epsilon\sqrt{t}}\right)\sim\mathcal{O}(\sqrt{T}).$
-B Proof of Strongly Convex and Smoothness

### v4

@@ Abstract
Online conformal prediction enables the runtime calibration of a pre-trained artificial intelligence model using feedback on its performance. Calibration is achieved through set predictions that are updated via online rules so as to ensure long-term coverage guarantees. While recent research has demonstrated the benefits of incorporating prior knowledge into the calibration process, this has come at the cost of replacing coverage guarantees with less tangible regret guarantees based on the quantile loss. This work introduces intermittent mirror online conformal prediction (IM-OCP), a novel runtime calibration framework that integrates prior knowledge, operates under potentially intermittent feedback, and features minimal memory complexity. IM-OCP guarantees long-term coverage and sub-linear regret, both of which hold deterministically for any given data sequence and in expectation with respect to the intermittent feedback.
@@ I Introduction
As summarized in Table I, existing online calibration schemes either cannot incorporate prior information or sacrifice long-term coverage guarantees (see Sec. II for details). This work addresses these limitations through the following key contributions:
$\bullet$ Intermittent mirror online conformal prediction (IM-OCP): We introduce IM-OCP, a novel online calibration framework that builds on online mirror descent (OMD) [21, 22, 23] by integrating prior knowledge and incorporating an importance weighting strategy [24] to handle intermittent feedback. IM-OCP exhibits constant memory usage and low-complexity updates.
$\bullet$ Regret and coverage guarantees: We demonstrate that IM-OCP guarantees long-term coverage and sublinear regret, both of which hold deterministically for any data sequence and in expectation over the intermittent feedback. While regret performance is basically inherited from OMD, establishing the coverage guarantees requires novel technical steps, including proving the boundedness of the IM-OCP iterates.
$\bullet$ Numerical validation: We showcase the merits of IM-OCP for the task of received signal strength indicator (RSSI)-based localization under intermittent feedback [25, 26, 27].
@@ II-A Online Calibration via Set Prediction
We consider the problem of online calibration in sequential decision-making problems. At each round $t$, a learner observes an input $X_{t}\in\mathcal{X}$ and aims to predict the corresponding label $Y_{t}\in\mathcal{Y}$. To this end, the predictor assigns a negatively oriented score $s(X_{t},Y)$ to each value $Y\in\mathcal{Y}$, which is assumed to be upper bounded by a constant $B<\infty$, i.e., $s(X,Y)\leq B$. Using predictor’s output, a prediction set $\mathcal{C}_{t}$ is produced that includes all labels $Y\in\mathcal{Y}$ with a score no larger than a threshold $r_{t}$, i.e., [10]
@@
In (2), the parameters $A$ and $\gamma$ are constants independent of $T$. By (2), the rate of miscoverage errors on an arbitrary sequence $\{(X_{t},Y_{t})\}_{t=1}^{T}$ converges to the target miscoverage level $\alpha$ as the number of steps $T$ tends to infinity.
@@ II-C Bayesian Adaptive Conformal Inference
The sublinear regret of IB-ACI can be easily established by leveraging standard results on the follow-the-regularized-leader [21, 22, 23] method.
@@ III-A1 Regularizer Encoding Prior Information
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$ for any fixed $\sigma>0$. In Appendix -A, we show that the function $R(\cdot)$ is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\mu$.
@@ III-A2 Threshold Update via OMD
Using the regularizer $R(r)$ in (9), IM-OCP updates the threshold at each time $t$ according to the OMD rule [21, 22, 23]. To describe it, define the mirror map induced by function $R(r)$ as $\mathcal{M}_{R}(\cdot)=\nabla R(\cdot)$, and the corresponding inverse as $\mathcal{M}_{R}^{-1}(\cdot)=(\nabla R)^{-1}(\cdot)$ The update rule of IM-OCP is then given as
@@
When the feedback $E_{t-1}$ is observed ($\text{obs}_{t-1}$$=$$1$), this update corresponds to a gradient step in the dual space defined by function $R(\cdot)$. This adjusts the threshold $r_{t}$ in a direction informed by both the observed feedback and the prior knowledge embedded in the mirror map $\mathcal{M}_{R}(\cdot)$. In contrast, when no feedback is observed ($\text{obs}_{t-1}$$=$00), the update in (10) reduces to the identity map, and the threshold remains unchanged.
Input: Miscoverage rate $\alpha$, and prior information $P(r)$
1 Initialization: $r_{1}=1-\alpha$
2 for $t=1,\dots,T$ do
3       Observe an input $X_{t}\in\mathcal{X}$
4       Specify the score function $s_{t}(X,Y)$
5       Return the prediction set $\mathcal{C}_{t}(X_{t},r_{t})$ via eq. (1)
6       Update the step size $\eta_{t}$ based on Corollary 1
7       if Observe Feedback $Y_{t}^{\star}$ ($\textnormal{obs}_{t}=1$) then
8             Update threshold $r_{t+1}$ via eq. (10)
9
10      else
11             Keep threshold $r_{t+1}=r_{t}$
12
13
Algorithm 1 IM-OCP
III-A3 Summary
The proposed IM-OCP is summarized in Algorithm 1. As shown, IM-OCP does not require storing past feedback values $Y_{t}^{\star}$, resulting in a constant memory. The inverse map $\mathcal{M}_{R}^{-1}(\cdot)$ can be evaluated offline using approximation methods such as neural network-based regression [28], entailing a minimal complexity and memory overhead.
@@ V Conclusion
-A Proof of Strongly Convex and Smoothness